# Possible Liquid-Nitrogen-Temperature Superconductivity Driven by Perpendicular Electric Field in the Single-Bilayer Film of La3Ni2O7 at Ambient Pressure

### v3

@@ Abstract
Recently, high-temperature superconductivity (HTSC) is found in the La3Ni2O7/SrLaAlO4 ultrathin film with critical temperature $T_{c}$ above the McMillan limit at ambient pressure (AP). It is eager to enhance $T_{c}$ of La3Ni2O7 at AP. We propose that a perpendicular electric field strongly enhances $T_{c}$ in the single-bilayer film of La3Ni2O7 at AP. Under electric field, the layer with lower potential energy will accept electrons flowing from the other layer to fill in the Ni-$3d_{x^{2}-y^{2}}$ orbitals, as the nearly half-filled Ni-$3d_{z^{2}}$ orbital cannot accommodate more electrons. With the enhancement of the filling fraction in the $3d_{x^{2}-y^{2}}$ orbitals in this layer, the interlayer $s$-wave pairing is suppressed, but the intralayer $d$-wave pairing in this layer is strongly enhanced. We numerically verify this idea and yield that an imposed voltage of about $0.1\sim 0.2$ volt between layers is enough to realize liquid-nitrogen-temperature HTSC in this single bilayer at AP. Our results appeal for experimental verification.
@@
The discovery of superconductivity (SC) with critical temperature $T_{c}$ above the boiling point of liquid nitrogen ($\approx 77\text{ K}$) in the pressurized La3Ni2O7 [Wang2023LNO, YuanHQ2023LNO, Wang2023LNOb, wang2023LNOpoly, wang2023la2prnio7, zhang2023pressure, zhou2023evidence, wang2024bulk, li2024pressure] has attracted great interests [Fukamachi2001, khasanov2024pressure, chen2024evidence, dan2024spin, chen2024electronic, Wang2022LNO, Kakoi2024, xie2024neutron, gupta2024anisotropic, ren2024resolving, feng2024unaltered, meng2024density, fan2024tunn, xu2024pressure, LI2024distinct, liu2024electronic, yashima2025microscopic, khasanov2025oxygen, yang2024orbital, wang2023structure, cui2023strain, Li2024ele, li2024distinguishing, zhou2024revealing, wang2025chemical, Chen2024poly, Dong2024vis, Li2024design, puphal2024unconven, zhu2024superconductivity, zhang2023superconductivity, huang2024signature, li2023trilayer, zhang2020intertwined, xu2024origin, du2024correlated, li2019superconductivity, lee2023linear, nomura2022superconductivity, gu2022superconductivity, sui2023rno, YaoDX2023, Dagotto2023, cao2023flat, zhang2023structural, huang2023impurity, geisler2023structural, rhodes2023structural, zhang2023la3ni2o6, yuan2023trilayer, li2024la3, geisler2024optical, li2017fermiology, wang2024non, chen2024tri, ZhangGM2023DMRG, Werner2023, shilenko2023correlated, WuWei2023charge, chen2023critical, ouyang2023hund, heier2023competing, wang2024electronic, botzel2024theory, WangQH2023, YangF2023, lechermann2023, Kuroki2023, HuJP2023, lu2023bilayertJ, oh2023type2, liao2023electron, qu2023bilayer, Yi_Feng2023, jiang2023high, zhang2023trends, qin2023high, tian2023correlation, jiang2023pressure, lu2023sc, kitamine2023, luo2023high, zhang2023strong, pan2023rno, sakakibara2023La4Ni3O10, lange2023mixedtj, yang2023strong, lange2023feshbach, kaneko2023pair, fan2023sc, wu2024deconfined, zhang2024prediction, zhang2024s, Yang2024effective, zhang2024electronic, yang2024decom, ryee2024quenched, Lu2024interplay, Ouyang2024absence, labollita2024, zhang2024emergent, Leonov2024Electronic, labollita2024assessing, ni2025spin, Yi2024nature, LaBollita2024electronic, jiang2024theory, chen2024non, zhang2024magnetic, lin2024magnetic, qin2024intertwined, Leonov2024Electronicc, liu2024growth, liu2024decoupling, tian2024spin, yin2025spmep, PhysRevB.111.104505, kaneko2025tj, Ji2025StrongCouplingLimit, Wang_2025, haque2025dft, shi2025theoretical, Ko2024signature, zhou2025ambient, liu2025superconductivity, yue2025correlated, li2025photoemission, bhatt2025resolving, shao2025band, shi2025effect, le2025landscape, Daoxin_Yao2025, wang2025electronic, huang2025spm, geisler2025electronic]. This discovery has sparked the exploration of high-temperature SC (HTSC) in Ruddlesden-Popper phase multilayer nickelates, resulting in the discovery of SC in the pressurized La4Ni3O10 [zhu2024superconductivity, zhang2023superconductivity, huang2024signature, li2023trilayer, zhang2020intertwined, xu2024origin, du2024correlated], which together with the previously synthesized infinite-layer nickelates Nd1-xSrxNiO2 [li2019superconductivity, lee2023linear, nomura2022superconductivity, gu2022superconductivity] have established a new family of SC other than cuprates and iron-based superconductors. However, the high pressure (HP) circumstance not only strongly hinders the experimental detection of the samples but also brings difficulties in the application of SC in industry. Very recently, the La3Ni2O7 ultrathin film with a few layers of unit cell grown on the SrLaAlO4 (SLAO) substrate has been grown by two different groups independently and SC with $T_{c}$ above the McMillan limit ($\approx$40 K) has been detected at ambient pressure (AP)[Ko2024signature, zhou2025ambient, liu2025superconductivity], allowing various experimental investigation of the pairing mechanism in this material, attracting a lot of interests [yue2025correlated, li2025photoemission, bhatt2025resolving, shao2025band, shi2025effect, le2025landscape, Daoxin_Yao2025, wang2025electronic, huang2025spm, geisler2025electronic]. It is now eager to enhance the $T_{c}$ of this material at AP. Here we propose a viable approach to realize $T_{c}$ above the boiling point of liquid nitrogen in the La3Ni2O7 single-bilayer film at AP.
@@
Presently, the pairing mechanism in the La3Ni2O7, either in the bulk material under HP [WangQH2023, YangF2023, lechermann2023, Kuroki2023, HuJP2023, lu2023bilayertJ, oh2023type2, liao2023electron, qu2023bilayer, Yi_Feng2023, jiang2023high, zhang2023trends, qin2023high, tian2023correlation, jiang2023pressure, lu2023sc, kitamine2023, luo2023high, zhang2023strong, pan2023rno, sakakibara2023La4Ni3O10, lange2023mixedtj, yang2023strong, lange2023feshbach, kaneko2023pair, fan2023sc, wu2024deconfined, zhang2024electronic, yang2024decom, ryee2024quenched, Lu2024interplay, Ouyang2024absence, yin2025spmep, PhysRevB.111.104505, kaneko2025tj, Ji2025StrongCouplingLimit] or in the ultrathin film at AP [yue2025correlated, shao2025band, le2025landscape, huang2025spm], is still under debate. Density-functional-theory (DFT) based first-principle calculations have suggested that the low-energy orbitals are mainly Ni-$3d_{z^{2}}$ and $3d_{x^{2}-y^{2}}$, which are nearly half- and quarter- filled [sui2023rno, YaoDX2023, Dagotto2023, cao2023flat, zhang2023structural, huang2023impurity, geisler2023structural, rhodes2023structural, zhang2023la3ni2o6, geisler2024optical]. Various experiments have revealed the strongly-correlated characteristic of the material [chen2024electronic, Kakoi2024, fan2024tunn, LI2024distinct, liu2024electronic, yang2024orbital, Li2024ele, li2024distinguishing]. Particularly, the optical study reveals significant reduction of the electron kinetic energy which places the system in the proximity of the Mott phase [liu2024electronic]; the angle-resolved photoemission spectroscopy uncovers strong band renormalization caused by electron correlation [yang2024orbital]; the linearly temperature-dependent resistivity suggests “strange-metal” behavior [YuanHQ2023LNO]. Therefore, we take a strong-coupling view of the system. Under the strong Hubbard repulsion, the nearly half-filled $3d_{z^{2}}$ electrons can almost be viewed as localized spins. Therefore, the main carrier of SC should be the $3d_{x^{2}-y^{2}}$ electrons, which subject to the in-plane superexchange interaction just mimic the 50% hole-doped cuprates. However, it is a problem how HTSC can emerge under such a high doping level. The key physics lies in the important role played by the $3d_{z^{2}}$ orbitals. Through strong interlayer perpendicular superexchange, the $3d_{z^{2}}$ electrons form interlayer pairing. The interlayer perpendicular superexchange or the interlayer pairing of the $3d_{z^{2}}$ electrons can be transmitted to the $3d_{x^{2}-y^{2}}$ electrons through the Hund’s rule [lu2023bilayertJ, oh2023type2, qu2023bilayer, tian2023correlation, zhang2023strong, pan2023rno, lange2023mixedtj, yang2023strong, lange2023feshbach, kaneko2023pair, wu2024deconfined, Lu2024interplay, Ouyang2024absence] or the nearest-neighbor (NN) hybridization [ZhangGM2023DMRG, Yi_Feng2023, qin2023high, tian2023correlation, luo2023high, kaneko2023pair, yang2024decom, Lu2024interplay, Ouyang2024absence] or both. Under such view, the role of pressure in enhancing the $T_{c}$ lies in the enhancement of the interlayer perpendicular superexchange, the inter-orbital hybridization, or both.
@@
In this work, we propose an alternative approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the ultrathin film of La3Ni2O7 at AP. Here we consider the thinnest limit, i.e. a single bilayer film of La3Ni2O7, and realize the goal by introducing charge transfer with a perpendicular electric field, which let the electrons flow from the high-energy layer to the low-energy layer, similar to the mechanism for the spontaneous charge transfer in oxide heterostructures [PhysRevB.95.205131, PhysRevLett.123.117201, PhysRevB.98.155103, PhysRevB.108.054441, mondal2023modulation, doi:10.1021/acs.nanolett.4c02087]. The external electric field based approach avoids introducing disorder as in chemical doping [jiang2021electronic] or exhibiting orbital selectivity based on symmetry [sohn2024observation], demonstrating exceptional performance in the field of twisted multilayer graphene materials [cao2020tunable, zhang2024observation, zhou2025gate]. We can impose a perpendicular electric field, say pointing upward, in this single bilayer, so that electrons from the top layer will flow to the bottom layer. These electrons will fill the $3d_{x^{2}-y^{2}}$ orbitals in the bottom layer as the nearly half-filled $3d_{z^{2}}$ orbitals there cannot accommodate more electrons. The enhancement of the bottom-layer $3d_{x^{2}-y^{2}}$ electron number will first suppress the interlayer $s$-wave SC due to mismatch of the electron numbers between the two layers, and then promptly lead to the intra-bottom-layer $d$-wave SC with strongly enhanced $T_{c}$. To test this idea, we have performed a combined simplified single orbital study and a comprehensive two orbital study, which consistently yield that a voltage of experimentally achievable levels (around $0.1\sim 0.2$ volt predicted by the mean-field calculations) between the two layers is enough to induce $d$-wave SC with $T_{c}$ above the boiling point of liquid nitrogen in the bottom layer. Intriguingly, the $d$-wave SC carried by the bottom layer $3d_{x^{2}-y^{2}}$ electrons coexists with the interlayer $s$-wave pseudo gap carried by the $3d_{z^{2}}$ electrons in the mixing ratio of $1:\mathrm{i}$, breaking time-reversal symmetry. Our proposal potentially provides a viable approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the single bilayer film of La3Ni2O7.
I Results
I.1 Consideration and a Simplified Study
@@
The La3Ni2O7 ultrathin film grown on the SLAO substrate form a bilayer square lattice [bhatt2025resolving, yue2025correlated]. As illustrated in Fig. 1 (a), the leading hopping integrals are the interlayer hopping of the $3d_{z^{2}}$ electrons $t_{\perp}$ and the intralayer NN hopping of the $3d_{x^{2}-y^{2}}$ electrons $t_{\parallel}$. Under strong Hubbard $U$, these hopping terms can induce the effective superexchange $J_{\perp}$ and $J_{\parallel}$ through $J\approx\frac{4t^{2}}{U}$. Under the Hund’s rule coupling $J_{H}$, the spins of the two orbitals are inclined to be parallel aligned, as illustrated in Fig. 1 (b), which partly transmits the interlayer perpendicular superexchange $J_{\perp}$ between the $3d_{z^{2}}$ orbitals to the $3d_{x^{2}-y^{2}}$ orbitals as $\tilde{J}_{\perp}=\alpha J_{\perp}$ with $\alpha\in(0,1)$ describing the efficiency of this process and related to the strength of Hund’s coupling $J_{H}$. In addition, there exists intralayer NN- hybridization $t_{xz}$ between the two orbitals. As shown in Fig. 2 (a, b), the nearly quarter-filled $3d_{x^{2}-y^{2}}$ electrons subject to $J_{\parallel}$ and $\tilde{J}_{\perp}$ form interlayer-dominant pairing [lu2023bilayertJ].
@@
Based on the above general consideration, we first conduct the following simplified model study including only the $3d_{x^{2}-y^{2}}$-orbital, with the $3d_{z^{2}}$ orbital only viewed as a source which tunes the total electron number. The following widely adopted single $3d_{x^{2}-y^{2}}$-orbital bilayer $t-J-J_{\perp}$ model [lu2023bilayertJ, oh2023type2, qu2023bilayer, zhang2023strong, pan2023rno, lange2023mixedtj, yang2023strong, wu2024deconfined] is adopted,
@@
The results indicate that (i) With the enhancement of the perpendicular electric field $\varepsilon$, and hence the transferred $d_{x^{2}-y^{2}}$-electron-doping level $\delta$, the pairing symmetry changes from interlayer $s$-wave to intra-bottom-layer $d$-wave (The criterion of the pairing symmetry is provided in METHODS); (ii) For all the transferred $d_{x^{2}-y^{2}}$-electron-doping levels $\delta$ tested, the enhancement of the perpendicular electric field $\varepsilon$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing; (iii) Under all the perpendicular electric field strengths $\varepsilon$ tested, the enhancement of the transferred $d_{x^{2}-y^{2}}$-electron-doping level $\delta$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing. From (ii) and (iii), it is clear that the enhancement of $\varepsilon$ and hence $\delta$ will significantly enhance the intra-bottom-layer pairing. These results are qualitatively consistent with those of the SBMF study. More results are given in the SI [SM].
Besides, we study the effect of interlayer Coulomb interaction. Our results show that the interlayer Coulomb interaction slightly promotes charge transfer between layers and the intra-bottom-layer pairing, while suppressing the interlayer pairing. See SI [SM] for details.
I.2 The comprehensive two-orbital study
The above simplified single-orbital study has drawbacks: We cannot determine the relationship between the electron-doping of the $d_{x^{2}-y^{2}}$ orbitals and the electric field. In the SBMF study, we have to assume the ratio between the changes of the filling fractions of the two top-layer $E_{g}$ orbitals. In addition, we do not know how the neglected $3d_{z^{2}}$ orbital degree of freedom affects the pairing nature. To settle these puzzles, we conduct a comprehensive two-orbital model [Lu2024interplay] to study with,

### v4

@@ Abstract
Abstract: Recently, high-temperature superconductivity (HTSC) is found in the La3Ni2O7/SrLaAlO4 ultrathin film with critical temperature $T_{c}$ above the McMillan limit at ambient pressure (AP). It is eager to enhance $T_{c}$ of La3Ni2O7 at AP. We propose that a perpendicular electric field strongly enhances $T_{c}$ in the single-bilayer film of La3Ni2O7 at AP. Under electric field, the layer with lower potential energy will accept electrons flowing from the other layer to fill in the Ni-$3d_{x^{2}-y^{2}}$ orbitals, as the nearly half-filled Ni-$3d_{z^{2}}$ orbital cannot accommodate more electrons. With the enhancement of the filling fraction in the $3d_{x^{2}-y^{2}}$ orbitals in this layer, the interlayer $s$-wave pairing is suppressed, but the intralayer $d$-wave pairing in this layer is strongly enhanced. We numerically verify this idea and yield that an imposed voltage of about $0.1\sim 0.2$ volt between layers is enough to realize liquid-nitrogen-temperature HTSC in this single bilayer at AP. Our results appeal for experimental verification.
I Introduction
@@
The discovery of superconductivity (SC) with critical temperature $T_{c}$ above the boiling point of liquid nitrogen ($\approx 77\text{ K}$) in the pressurized La3Ni2O7 [111, 150, 30, 115, 117, 149, 161, 120, 65] has attracted great interests [21, 44, 9, 158, 10, 78, 1, 129, 27, 99, 20, 85, 18, 130, 70, 79, 138, 45, 133, 119, 15, 71, 66, 160, 116, 11, 16, 62, 96, 163, 148, 34, 69, 146, 131, 17, 61, 56, 88, 25, 110, 83, 151, 5, 153, 32, 23, 100, 155, 141, 64, 22, 63, 118, 8, 105, 14, 108, 126, 12, 91, 29, 121, 4, 135, 75, 55, 102, 26, 80, 89, 72, 98, 136, 39, 154, 97, 113, 40, 82, 46, 84, 145, 92, 103, 52, 132, 53, 43, 19, 127, 152, 147, 134, 156, 137, 101, 81, 90, 51, 143, 59, 50, 87, 139, 49, 38, 13, 157, 73, 36, 60, 76, 77, 112, 140, 128, 42, 35, 122, 28, 107, 47, 159, 74, 142, 67, 3, 104, 106, 54, 31, 114, 33, 24]. This discovery has sparked the exploration of high-temperature SC (HTSC) in Ruddlesden-Popper phase multilayer nickelates, resulting in the discovery of SC in the pressurized La4Ni3O10 [163, 148, 34, 69, 146, 131, 17], which together with the previously synthesized infinite-layer nickelates Nd1-xSrxNiO2 [61, 56, 88, 25] have established a new family of SC other than cuprates and iron-based superconductors. However, the high pressure (HP) circumstance not only strongly hinders the experimental detection of the samples but also brings difficulties in the application of SC in industry. Very recently, the La3Ni2O7 ultrathin film with a few layers of unit cell grown on the SrLaAlO4 (SLAO) substrate has been grown by two different groups independently and SC with $T_{c}$ above the McMillan limit ($\approx$40 K) has been detected at ambient pressure (AP)[47, 159, 74], allowing various experimental investigation of the pairing mechanism in this material, attracting a lot of interests [142, 67, 3, 104, 106, 54, 31, 114, 33, 24]. It is now eager to enhance the $T_{c}$ of this material at AP. Here we propose a viable approach to realize $T_{c}$ above the boiling point of liquid nitrogen in the La3Ni2O7 single-bilayer film at AP.
@@
Presently, the pairing mechanism in the La3Ni2O7, either in the bulk material under HP [135, 75, 55, 102, 26, 80, 89, 72, 98, 136, 39, 154, 97, 113, 40, 82, 46, 84, 145, 92, 103, 52, 132, 53, 43, 19, 127, 156, 137, 101, 81, 90, 140, 128, 42, 35] or in the ultrathin film at AP [142, 104, 54, 33], is still under debate. Density-functional-theory (DFT) based first-principle calculations have suggested that the low-energy orbitals are mainly Ni-$3d_{z^{2}}$ and $3d_{x^{2}-y^{2}}$, which are nearly half- and quarter- filled [110, 83, 151, 5, 153, 32, 23, 100, 155, 22]. Various experiments have revealed the strongly-correlated characteristic of the material [10, 1, 18, 70, 79, 133, 71, 66]. Particularly, the optical study reveals significant reduction of the electron kinetic energy which places the system in the proximity of the Mott phase [79]; the angle-resolved photoemission spectroscopy uncovers strong band renormalization caused by electron correlation [133]; the linearly temperature-dependent resistivity suggests “strange-metal” behavior [150]. Therefore, we take a strong-coupling view of the system. Under the strong Hubbard repulsion, the nearly half-filled $3d_{z^{2}}$ electrons can almost be viewed as localized spins. Therefore, the main carrier of SC should be the $3d_{x^{2}-y^{2}}$ electrons, which subject to the in-plane superexchange interaction just mimic the 50% hole-doped cuprates. However, it is a problem how HTSC can emerge under such a high doping level. The key physics lies in the important role played by the $3d_{z^{2}}$ orbitals. Through strong interlayer perpendicular superexchange, the $3d_{z^{2}}$ electrons form interlayer pairing. The interlayer perpendicular superexchange or the interlayer pairing of the $3d_{z^{2}}$ electrons can be transmitted to the $3d_{x^{2}-y^{2}}$ electrons through the Hund’s rule [80, 89, 98, 113, 145, 92, 52, 132, 53, 43, 127, 81, 90] or the nearest-neighbor (NN) hybridization [105, 136, 97, 113, 84, 43, 137, 81, 90] or both. Under such view, the role of pressure in enhancing the $T_{c}$ lies in the enhancement of the interlayer perpendicular superexchange, the inter-orbital hybridization, or both.
@@
In this work, we propose an alternative approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the ultrathin film of La3Ni2O7 at AP. Here we consider the thinnest limit, i.e. a single bilayer film of La3Ni2O7, and realize the goal by introducing charge transfer with a perpendicular electric field, which let the electrons flow from the high-energy layer to the low-energy layer, similar to the mechanism for the spontaneous charge transfer in oxide heterostructures [95, 57, 7, 93, 86, 94]. The external electric field based approach avoids introducing disorder as in chemical doping [37] or exhibiting orbital selectivity based on symmetry [109], demonstrating exceptional performance in the field of twisted multilayer graphene materials [6, 144, 162]. We can impose a perpendicular electric field, say pointing upward, in this single bilayer, so that electrons from the top layer will flow to the bottom layer. These electrons will fill the $3d_{x^{2}-y^{2}}$ orbitals in the bottom layer as the nearly half-filled $3d_{z^{2}}$ orbitals there cannot accommodate more electrons. The enhancement of the bottom-layer $3d_{x^{2}-y^{2}}$ electron number will first suppress the interlayer $s$-wave SC due to mismatch of the electron numbers between the two layers, similar with the case in which an imposed Zeeman field acting on the spin leads to mismatch of the electron numbers between the two spin species and thus suppresses singlet pairing, and then promptly lead to the intra-bottom-layer $d$-wave SC with strongly enhanced $T_{c}$. To test this idea, we have performed a combined simplified single orbital study and a comprehensive two orbital study, which consistently yield that a voltage of experimentally achievable levels (around $0.1\sim 0.2$ volt predicted by the mean-field calculations) between the two layers is enough to induce $d$-wave SC with $T_{c}$ above the boiling point of liquid nitrogen in the bottom layer. Intriguingly, the $d$-wave SC carried by the bottom layer $3d_{x^{2}-y^{2}}$ electrons coexists with the interlayer $s$-wave pseudo gap carried by the $3d_{z^{2}}$ electrons in the mixing ratio of $1:\mathrm{i}$, breaking time-reversal symmetry. Our proposal potentially provides a viable approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the single bilayer film of La3Ni2O7.
II Results
II.1 Consideration and a Simplified Study
@@
The La3Ni2O7 ultrathin film grown on the SLAO substrate form a bilayer square lattice [3, 142]. As illustrated in Fig. 1 (a), the leading hopping integrals are the interlayer hopping of the $3d_{z^{2}}$ electrons $t_{\perp}$ and the intralayer NN hopping of the $3d_{x^{2}-y^{2}}$ electrons $t_{\parallel}$. Under strong Hubbard $U$, these hopping terms can induce the effective superexchange $J_{\perp}$ and $J_{\parallel}$ through $J\approx\frac{4t^{2}}{U}$. Under the Hund’s rule coupling $J_{H}$, the spins of the two orbitals are inclined to be parallel aligned, as illustrated in Fig. 1 (b), which partly transmits the interlayer perpendicular superexchange $J_{\perp}$ between the $3d_{z^{2}}$ orbitals to the $3d_{x^{2}-y^{2}}$ orbitals as $\tilde{J}_{\perp}=\alpha J_{\perp}$ with $\alpha\in(0,1)$ describing the efficiency of this process and related to the strength of Hund’s coupling $J_{H}$. In addition, there exists intralayer NN- hybridization $t_{xz}$ between the two orbitals. As shown in Fig. 2 (a, b), the nearly quarter-filled $3d_{x^{2}-y^{2}}$ electrons subject to $J_{\parallel}$ and $\tilde{J}_{\perp}$ form interlayer-dominant pairing [80].
@@
Based on the above general consideration, we first conduct the following simplified model study including only the $3d_{x^{2}-y^{2}}$-orbital, with the $3d_{z^{2}}$ orbital only viewed as a source which tunes the total electron number. The following widely adopted single $3d_{x^{2}-y^{2}}$-orbital bilayer $t-J-J_{\perp}$ model [80, 89, 98, 145, 92, 52, 132, 127] is adopted,
@@
The results indicate that (i) With the enhancement of the perpendicular electric field $\varepsilon$, and hence the transferred $d_{x^{2}-y^{2}}$-electron-doping level $\delta$, the pairing symmetry changes from interlayer $s$-wave to intra-bottom-layer $d$-wave (The criterion of the pairing symmetry is provided in METHODS); (ii) For all the transferred $d_{x^{2}-y^{2}}$-electron-doping levels $\delta$ tested, the enhancement of the perpendicular electric field $\varepsilon$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing; (iii) Under all the perpendicular electric field strengths $\varepsilon$ tested, the enhancement of the transferred $d_{x^{2}-y^{2}}$-electron-doping level $\delta$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing. From (ii) and (iii), it is clear that the enhancement of $\varepsilon$ and hence $\delta$ will significantly enhance the intra-bottom-layer pairing. These results are qualitatively consistent with those of the SBMF study. More results are given in the SI [2].
Besides, we study the effect of interlayer Coulomb interaction. Our results show that the interlayer Coulomb interaction slightly promotes charge transfer between layers and the intra-bottom-layer pairing, while suppressing the interlayer pairing. See SI [2] for details.
II.2 The comprehensive two-orbital study
The above simplified single-orbital study has drawbacks: We cannot determine the relationship between the electron-doping of the $d_{x^{2}-y^{2}}$ orbitals and the electric field. In the SBMF study, we have to assume the ratio between the changes of the filling fractions of the two top-layer $E_{g}$ orbitals. In addition, we do not know how the neglected $3d_{z^{2}}$ orbital degree of freedom affects the pairing nature. To settle these puzzles, we conduct a comprehensive two-orbital model [81] to study with,